OpenStudy (anonymous):

Solve The equation. 1 1 1 - + - = - x 2x 6 x

OpenStudy (anonymous):

thats not an equation

OpenStudy (anonymous):

it looks like a table?

OpenStudy (shamil98):

\[\frac{ 1 }{ x } + \frac{ 1 }{ 2x } = \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

ohhhh i see i see i see

OpenStudy (anonymous):

Yeah that's what it's supposed to be. ^^

OpenStudy (anonymous):

haha lol well

OpenStudy (anonymous):

I thought I could multiply it all by 6x,but it didn't work???

OpenStudy (shamil98):

find a common denominator the fractions with the variables and then add. so 1/2 so.. \[\frac{ 2 }{ 2x} + \frac{ 1 }{ 2x } = \frac{ 1 }{ 6 } \] \[\frac{ 3 }{ 2x } = \frac{ 1 }{ 6 }\] cross multiply. \[18 = 2x\] x = 9

OpenStudy (anonymous):

so let's say it's \[\frac{ x }{ x+2}+ \frac{ 4 }{ x-2 } =1\] would I multiply both fraction so you get \[x^{2}-4\] as the denominator?